Sol Summary – May 20th

Crew 281 Sol Summary Report 20 May 2023
Sol: 6
Summary Title:
Author’s name: Ritupriya Patil (XO) and Megan Kane (Commander)
Mission Status: Nominal
Sol Activity Summary:
First EVA was at the Kissing Camel Ridge. The crew of 3 drove Opportunity and Curiosity rovers and reached the site at 0755 hrs. The first objective was to capture scanline measurements for the boulders Ana Pires (Crew Scientist) approved on the East ridge. KC Shasteen helped with the Equotip measurements while Ana gathered the Schmidt hammer data. After finishing 5 station long scanline. Ritupriya Patil (XO, EVA Lead) set up to deploy the drone. Detailed vertical images and videos of the scanline were captured to help confirm the weathering grade, discontinuities of the rocks. There were also 4 translational drone passes along the scanline that captured the details of the boulders under geophysical analysis. These images and videos will be used for georeferencing the area around the scanline and for further assessment of the geological region.
For the second EVA the crew of 2 erected ham radio antenna to the northwest of the Hab. A coaxial cable was retrieved from where it had been passed through the tunnel wall. The cable length dictated where the station could be installed. It took 3 attempts to find a location where the ground spike could be deployed successfully. Once deployed the ground wire and the mini were attached. This was followed by the antenna and lastly by the coaxial cable. Grounding pole could not be deployed, and the wire was affixed to a rock. The hazard cones were placed, and the crew returned to the hab.
After the ham radio station set up Rachel Jones (HSO) jumped on the radio and established contact with Dan from Govetown, Gorgia, N4MI. This ham radio testing was approved during the Karman line discussion.
In the Greenhab, Megan (Commander) watered the plants and filled the cooler.
The science dome was quiet today, cacao seeds still doing well.
In the RAM, Ritu (XO) worked on her payload carrying harness for tomorrow’s drone flight test. The objective is to test short distance payload drop off for a simulated medical emergency.
For lunch Ritu (XO) made tomato sauce based pasta with mushrooms. And dinner was stir-fried veggies in soy sauce served with rice, made by KC. After that we buried our heads in report writing and mid-mission research material.
Look Ahead Plan: Tomorrow we will do an EVA for 3 around the hab to run second test on the Pegasus scoop (rock sampling device) and do a drone payload test and drill for minimum payload weight and distance. Likely do some work around the ham radio antenna.
Anomalies in work: None
Weather: Nominal
Crew Physical Status: Nominal
EVA: Completed 2 today

Reports to be filed:
Greenhab Report
Ops Report
Journalist Report
Photo of the Day

Support Requested:
Mission support is requested in case of storms, if the antenna needs to come down or be retrieved.

Copyright © The Mars Society. All rights reserved. | Main Site